# Modeling bacterial flow field with regularized singularities

### v1

@@ Abstract
The above needs to understand the near-field hydrodynamics of bacteria, have motivated sustained theoretical attempts [10, 38, 1], where the flow regularization method is taken as a promising candidate. Regularization, a mathematical technique that replaces singular solutions in hydrodynamic computations with smooth analytical functions, prevents the flow field from diverging near the singularity points and facilitates approachable numerical computations [31]. For example, a regularized Stokeslet is usually implemented by replacing the point force, specifically a Dirac delta function, with a smooth function (blob function) of certain geometry and characteristic length scale, and it has been widely utilised for modeling ciliary and flagellar flows [8, 11, 30, 37]. The flow field generated by a swimming bacterium is usually more complicated, due to its irregular geometry and complex swimming mechanism, despite some initial but meaningful trials, such as generating the flow field by combining a regularized stresslet and rotlet dipole [17]. However, most of the current regularization methods use isotropic blob functions, which are inherently unable to account for the anisotropy of the bacterial body and to capture the accurate near-body flow.
@@
Figure 2: (a) Entire flow field (ARS + ISRD) given by the model in the free space. Colorbar indicates the magnitude of the in-plane flow velocity $|u|=\sqrt{u_{x}^{2}+u_{y}^{2}}$. Flow field of a stresslet of the same strength as the ARS is presented in the inset. ARS: anisotropically regularized stresslet, IRSD: isotropically regularized source dipole. (b) The head and side flow profile $|u_{x}(x,0,0)|$ (red) and $|u_{y}(0,y,0)|$ (blue). Experimental measurements are represented by dots, numerical simulations by diamonds. Theoretical predictions inside and outside the BEM cell body surface are shown as dashed and solid lines, respectively.
Consider a bacterium whose body center is located at the origin and the body axis is oriented along +x+x, as shown in Fig. 1, and then the flow field generated by the bacterium can be modelled as a superposition of the flow fields generated by an anisotropically regularized stresslet (ARS) and an isotropically regularized source dipole (IRSD).
@@
where the distance $r_{\parallel}=|\bm{q}\cdot\bm{r}|$ and $r_{\perp}=|\bm{r}-r_{\parallel}\bm{q}|$, and $\varepsilon_{\parallel}$ and $\varepsilon_{\perp}$ are two regularization parameters along and perpendicular to the body axis oriented in the direction $\bm{q}$, respectively. Note that the regularization parameter reflect the approximate distance at which the flow velocity reaches its extremum along the corresponding direction.
@@
Then, the entire flow field in the laboratory frame, is given by the superposition of the above two flow field generated by the ARS and IRSD. For simplicity, we set $\bm{q}$ as +x+x direction. Since the center of the cell body is located at the origin and flagellar bundle extends to $x<0$, ARS center is placed at $\bm{r}_{A}=(-l,0,0)$, as shown in Fig. 1. Because IRSD enforces no-slip condition over the cell body which is symmetric around the origin, its center coincides with the geometric center of the cell body. Therefore, the entire flow field is given by,
@@
Note that, since the regularization does not alter the far-field behavior of singularity solutions, the strength of both the stresslet $\bm{D}$ and the source dipole $\bm{d}$ are defined in the same fashion as those in the case without regularization.
@@
In the free space, both the experimental and BEM results are accurately captured by the theoretical predictions using the parameter set $\{D=1\ \text{pN}\cdot\mu\text{m},l=1.6\ \mu\text{m},\varepsilon_{\parallel}=1~\mu\text{m},\varepsilon_{\perp}=1.36\ \mu\text{m},d=140\ \mu\text{m}^{4}/\text{s},\lambda=0.7\ \mu\text{m}\}$. The value of the stresslet strength, $D$, is very close to the experimentally measured stresslet strength of E. coli [15], and the size, $l$, is comparable to the semi-major axis length of the cell body of E. coli [23, 5]. The fluid viscosity, $\mu=1\times 10^{-3}~\text{Pa}\cdot\text{s}$ is used in calculations.
The pattern of the flow field in the $xy$-plane, calculated by the model, is provided in Fig. 2(a). In the far field limit, the flow velocity decays in the form of $\propto r^{-2}$ as shown in in Fig. 2(b), same as that by a singular stresslet. This is due to the fast decay of the flow field given by IRSD, which is in the form of $\propto r^{-3}$, and the effect of IRSD then becomes negligble in the far field limit. In the near field limit, the model captures two main features observed in experiments and BEM simulations: (a) the head flow profile, $u_{x}(x,0,0)$, decays monotonically and the swimming speed is close to that obtained in BEM simulations, 11.8 $\mu m$/s, which is mainly due to the IRSD; (b) the side flow profile, $u_{y}(0,y,0)$, has an initial increase followed by a decrease in the form of $\propto r^{-2}$, and the flow speed approaches zero at the surface $y=1~\mu\text{m}$, due to the flow inversion induced by the stresslet center offset $l$.
@@
The flow field parallel to the surface ($xy$-plane) and perpendicular to the surface ($xz$-plane) are shown in Figs. 3 (a) and (b), respectively. The flow profiles along the head and side directions are shown in Fig. 3(c). The flow field in the near field limit behaves similarly as in the free space, whereas the velocity decaying in the far field limit follows the form of $\propto r^{-3}$ due to the no-slip wall. Note that both the theoretical results and BEM simulations show that the extremum of the side flow profile is lower than that in the free space, whereas the experimental measurements show the opposite trend: the extremum of the side flow profile being higher in the presence of the wall. This discrepancy can originate from a spiral flow field, which is usually described by a pair of opposite rotlets [17, 24, 21](see Supplementary Material).
@@
We develop an analytical model for describing the flow field generated by a swimming bacterium, in both the far field and the near field limit, based on the combination of regularized singularities. The theoretical results are validated by comparisons with experiments and our BEM simulations. The anisotropy introduced in the model nicely captures the geometry of a bacterium, and thus the flow field. The approach provides a complementary view to the multipolar expansion of microswimmer flow field [7, 22] and can boost further investigations in flow fields generated by other microorganisms [13, 34, 33]. Moreover, the flow field predicted by the model can be utilised to describe the hydrodynamic interactions of bacteria in dense suspensions, which is important in determining the collective dynamics of bacteria, e.g., bacterial turbulence.

### v2

@@ Abstract
The above needs to understand the near-field hydrodynamics of bacteria, have motivated sustained theoretical attempts [10, 43, 1], where the flow regularization method is taken as a promising candidate. Regularization, a mathematical technique that replaces singular solutions in hydrodynamic computations with smooth analytical functions, prevents the flow field from diverging near the singularity points and facilitates approachable numerical computations [35]. For example, a regularized Stokeslet is usually implemented by replacing the point force, specifically a Dirac delta function, with a smooth function (blob function) of certain geometry and characteristic length scale, and it has been widely utilised for modeling ciliary and flagellar flows [8, 11, 34, 42]. For the flow field generated by a microswimmer, the description based on regularized singularities remains underexplored. This is partly because the flow field generated by a swimming microorganism—even one as simple as a bacterium—is typically complex, owing to its irregular geometry and intricate swimming mechanism. Despite some initial but meaningful efforts, such as representing the flow field by combining a regularized stresslet and rotlet dipole [19], until now, these theoretical attempts do not quantitatively capture the experimentally observed near-field flow structure, which is essential for determining bacterial hydrodynamic interaction in dense suspension. Moreover, most of the current regularization methods use isotropic blob functions, which are inherently unable to account for the anisotropy of the bacterial body and to capture the accurate near-body flow.
@@
Figure 2: (a) Entire flow field (ARS + ISRD) given by the model in the free space. Colorbar indicates the magnitude of the in-plane flow velocity $|u|=\sqrt{u_{x}^{2}+u_{y}^{2}}$. Flow field of a stresslet of the same strength as the ARS is presented in the inset. ARS: anisotropically regularized stresslet, IRSD: isotropically regularized source dipole. (b) The head and side flow profile $|u_{x}(x,0,0)|$ (red) and $|u_{y}(0,y,0)|$ (blue). Experimental measurements are represented by dots, numerical simulations by diamonds. Theoretical predictions inside and outside the BEM cell body surface are shown as dashed and solid lines, respectively. The fluid viscosity, $\mu=1\times 10^{-3}~\text{Pa}\cdot\text{s}$ is used in calculations throughout this work.
Consider a bacterium whose body center is located at the origin and the body axis is oriented along +x+x, as shown in Fig. 1, and then the flow field generated by the bacterium can be modelled as a superposition of the flow fields generated by an anisotropically regularized stresslet (ARS) and an isotropically regularized source dipole (IRSD). Note the anisotropy introduced in ARS can help to characterize the flow field induced by non-spherical shape of the swimmer.
@@
where the distance $r_{\parallel}=|\bm{q}\cdot\bm{r}|$ and $r_{\perp}=|\bm{r}-r_{\parallel}\bm{q}|$, and $\varepsilon_{\parallel}$ and $\varepsilon_{\perp}$ are two regularization parameters along and perpendicular to the body axis oriented in the direction $\bm{q}$, respectively. For a regularized stresslet (located at the origin), the regularization parameters generally correspond to the distance at which the flow velocities $u_{x}(x,0,0)$ and $u_{y}(0,y,0)$ reach their extrema along $x$ and $y$ axis, respectively (see more details in Supplementary Material).
@@
Then, the entire flow field in the laboratory frame, is given by the superposition of the above two flow field generated by the ARS and IRSD. For simplicity, we set $\bm{q}$ as +x+x direction. Since the center of the cell body is located at the origin and flagellar bundle extends to $x<0$, ARS center is placed at $\bm{r}_{A}=(-l,0,0)$, as shown in Fig. 1. Because IRSD enforces no-slip condition at the two ends of the long axis of the cell body, which is symmetric regarding the origin, its center coincides with the geometric center of the cell body. Therefore, the entire flow field is given by,
@@
Note that, since the regularization does not alter the far-field behavior of singularity solutions, the strength of both the stresslet $\bm{D}$ and the source dipole $\bm{d}$ are defined in the same fashion as those in the case without regularization. For readers familiar with the squirmer model [18, 20, 15, 27], a higher-order multipole summation is usually needed for accurately capturing the flow field near the cell body, which is computationally heavy.
@@
[…]
The pattern of the flow field in the $xy$-plane, calculated by the model, is provided in Fig. 2(a). In the far field limit, the flow velocity decays in the form of $\propto r^{-2}$ as shown in in Fig. 2(b), same as that by a singular stresslet. This is due to the fast decay of the flow field given by IRSD, which is in the form of $\propto r^{-3}$, and the effect of IRSD then becomes negligble in the far field limit. In the near field limit, the model captures two main features observed in experiments and BEM simulations: (a) the head flow profile, $u_{x}(x,0,0)$, decays monotonically and the swimming speed is close to that obtained in BEM simulations, 11.8 $\mu m$/s; (b) the side flow profile, $u_{y}(0,y,0)$, has an initial increase followed by a decrease in the form of $\propto r^{-2}$, and the flow speed approaches zero at the surface $y=1~\mu\text{m}$, due to the flow inversion induced by the stresslet center offset $l$.
@@
The flow field parallel to the surface ($xy$-plane) and perpendicular to the surface ($xz$-plane) are shown in Figs. 3 (a) and (b), respectively. The flow profiles along the head and side directions are shown in Fig. 3(c). The flow field in the near field limit behaves similarly as in the free space, whereas the velocity decaying in the far field limit follows the form of $\propto r^{-3}$ due to the no-slip wall. Note that both the theoretical results and BEM simulations show that the extremum of the side flow profile is lower than that in the free space, whereas the experimental measurements show the opposite trend: the extremum of the side flow profile being higher in the presence of the wall. This discrepancy can originate from a spiral flow field, which is usually described by a pair of opposite rotlets [19, 28, 24]. Note that when comparing $u_{y}$ in Fig. 3, we have removed the contribution of the rotlet dipole by taking $u_{y}\left(0,y,0\right)=\left[u_{y}^{\prime}\left(0,y,0\right)-u_{y}^{\prime}\left(0,-y,0\right)\right]/2$ where $u_{y}^{\prime}$ denotes the raw flow field in BEM (see Supplementary Material).
@@
We develop an analytical model for describing the flow field generated by a swimming bacterium, in both the far field and the near field limit, based on the combination of regularized singularities. The theoretical results are validated by comparisons with experiments and our BEM simulations. The anisotropy introduced in the model nicely captures the geometry of a bacterium, and thus the flow field. The approach provides a complementary view to the multipole expansion of microswimmer flow field [7, 25] and can boost further investigations in flow fields generated by other microorganisms [13, 37, 36]. Moreover, the flow field predicted by the model can be utilised to describe the hydrodynamic interactions (forces and torques) between bacteria in dense suspensions, which is important in determining the collective dynamics of bacteria, e.g., bacterial turbulence.